A FIELD IS 32 FEET LONG SIX YARDS WIDE. WHAT IS THE PERMITER IN FEET.

Answer by checkley79(3341)

P=2L+2W
P=2*32+2(6*3)
P=64+2*18
P=64+36
P=100 FEET ANS.